OpenStudy (anonymous):

In a triangle, the measure of the first angle is 24 degrees more than the measure of the second angle. The measure of the third angle is four times the measure of the second angle. If the total number of degrees in a triangle is 180, what is the measure of the largest angle? 26 degrees 50 degrees 104 degrees 136 degrees

OpenStudy (alexandervonhumboldt2):

a=b+24 c=4*(b) a+b+c=180

OpenStudy (agl202):

Ya, thats what I got!!! @AlexandervonHumboldt2